Use the sqare root property to solve.
1.) y2=200
2.) (x+8)^2=25

Answer by Nate(3500)

1.) y^2 = 200
y = +- sqrt(200)
y = +- sqrt(2*2*2*5*5)
y = +- 10sqrt(2)
2.) (x + 8)^2 = 25
(x + 8)^2 = 25
x + 8 = +- 5
x = -8 +- 5
x = -13 and x = -3